OpenStudy (anonymous):

The equation of the line containing one side of a parallelogram is 3x + 2y = 8. The opposite side contains the point (0, -7). what is the equation of the line that contains the opposite side?

OpenStudy (anonymous):

the slope of the line you mentioned is -3/2 so that means that the opposite side must have an equation whose slope = -3/2 it also passes through the point (0,-7) so use the following general equation to find this y - y1 = m(x - x1) where m = -3/2 , x1 = 0 and y1 = -7

OpenStudy (anonymous):

thank you darling (: your the best (: